# Adapting Conformal Prediction to Distribution Shifts Without Labels

### v1

@@ Abstract
Conformal prediction (CP) enables machine learning models to output prediction sets with guaranteed coverage rate, assuming exchangeable data. Unfortunately, the exchangeability assumption is frequently violated due to distribution shifts in practice, and the challenge is often compounded by the lack of ground truth labels at test time. Focusing on classification in this paper, our goal is to improve the quality of CP-generated prediction sets using only unlabeled data from the test domain. This is achieved by two new methods called ECP and EACP, that adjust the score function in CP according to the base model’s uncertainty on the unlabeled test data. Through extensive experiments on a number of large-scale datasets and neural network architectures, we show that our methods provide consistent improvement over existing baselines and nearly match the performance of supervised algorithms.
@@ 1 Introduction
Advances in deep learning are fundamentally changing the autonomous decision making pipeline. While most works have focused on accurate point predictions, calibrating the uncertainty of the model is arguably as important. Taking autonomous driving as an example: if a detection model predicts the existence of an obstacle, it would be reasonable to take different maneuvering strategies depending on the confidence of the prediction. But is that reliable? In a possible failure mode, the model could report $60\%$ (resp. $99\%$) confidence, but the probability of an obstacle actually showing up is $99\%$ (resp. $60\%$). Such discrepancy between the model’s own uncertainty evaluation and the ground truth probability (or post-hoc frequency) is actually very common (Guo et al., 2017; Liang et al., 2023), and can significantly compromise the safety in downstream decision making.
@@
Conformal Prediction (CP; Vovk et al., 2005) is a simple post-hoc and distribution-free method that addresses this problem. Given a fixed black-box machine learning model (called the base model) and a covariate $x_{\mathrm{test}}$, the goal of CP is to generate a small prediction set $\mathcal{C}_{\mathrm{test}}$ that contains (or, covers) the unknown ground truth label $y_{\mathrm{test}}$ with a pre-specified probability. Crucially, CP relies on the assumption that the distribution of the data stream is exchangeable (a weaker variant of i.i.d.), which allows the fairly straightforward inference of $y_{\mathrm{test}}$ from $x_{\mathrm{test}}$ and the base model’s performance on a pre-collected calibration dataset. Note that the target label $y_{\mathrm{test}}$ does not need to be revealed after the set prediction is made: exchangeability together with a large enough calibration dataset is sufficient to ensure the desirable coverage probability. This is particularly important for autonomous decision making, where real-time data annotation is expensive or even infeasible.
@@
However, real-world data streams are usually corrupted by all sorts of distribution shifts, violating the exchangeability assumption. Even when the data stream itself is exchangeable, we often want to continually update the base model rather than keeping it fixed, and this can be effectively understood as a distribution shift in the CP context. In such cases, simply applying exchangeability-based CP methods could fail miserably (Tibshirani et al., 2019; Bhatnagar et al., 2023a; Kasa and Taylor, 2023). Therefore, making CP compatible with distribution shifts has become a focal point of recent works.
@@
A number of solutions have been proposed, but the key challenge still remains. For example, Gibbs and Candes (2021) formulated the connection between CP and Online Convex Optimization (OCO; Zinkevich, 2003), and the latter is able to handle arbitrarily distribution-shifted environments. The weakness is that ground truth labels are now required at test time, as opposed to the standard CP procedure. On the other direction, there are CP methods that combat distribution shifts without test time labels (Tibshirani et al., 2019; Roth, 2022; Barber et al., 2023; Cauchois et al., 2024), but they assume the distribution shifts are “easy”, such that even without labels, we can still rigorously infer the test distribution to a certain extent using the labeled calibration dataset. Overall, it appears that handling both difficulties – distribution shifts and the lack of test time labels – is a formidable but important challenge remaining in the CP literature.
@@
In this paper, we focus on classification, and address the above problem by exploiting the uncertainty evaluation of the base model itself. Although such a quantity is not necessarily calibrated in a strict sense, it has been empirically shown to strongly correlate with the magnitude of distribution shifts (Wang et al., 2021), thus providing a valuable way to probe the test distribution without label access. Under this high level idea, we make the following contributions:
@@
  • First, we propose a new method named ECP (Entropy scaled Conformal Prediction), which scales the score function in CP by an “entropy quantile” [^1] Enforced to be larger than 1 to ensure that the prediction sets do not become smaller on in-distribution data. of the base model on the test distribution.
@@
    In the CP jargon, the score function is determined by the fixed base machine learning model, and assigns each candidate label a “propensity score” after observing the covariate $x_{\mathrm{test}}$. Then, the CP prediction set $\mathcal{C}_{\mathrm{test}}$ simply includes all the labels whose score is above a certain threshold. [^2] Since we mainly focus on classification, the score functions are assumed to be positively oriented (Sadinle et al., 2019). By scaling up the score function, ECP makes the prediction sets larger, which naturally corresponds to the intuition that the uncertainty of the base model should be inflated under distribution shifts. Moreover, the amount of such inflation is correlated with the magnitude of the distribution shift, through the use of the entropy quantile.
@@
  • Second, we refine the above using techniques from unsupervised Test Time Adaptation (TTA) (Niu et al., 2022), and the resulting method is named as EACP (Entropy base-Adapted Conformal Prediction). The key idea is that while ECP keeps the base model fixed during testing, we may concurrently update it using entropy minimization (Grandvalet and Bengio, 2004; Wang et al., 2021) – a widely adopted idea in unsupervised TTA, alongside the aforementioned entropy scaling. This “adaptively” reduces the scaling effect that ECP applies on the score function, thus making the prediction sets of ECP smaller.
@@
  • Finally, despite the loss of statistical guarantees, we evaluate the proposed methods on a wide range of large-scale datasets under distribution shifts, as well as different neural network architectures. We find that exchangeability-based CP (with and without TTA on the base model) consistently leads to lower-than-specified coverage frequency, but our methods can effectively mitigate this under-coverage issue while keeping the sizes of the prediction sets moderate. Furthermore, our methods also significantly improve the prediction sets generated by the base model itself (without CP). It shows that by bridging the CP procedure (which is statistically sound) and the base model’s own uncertainty evaluation (which is often informative), our methods enjoy the practical benefit from both worlds.
2 Related works
CP in decision making
Our interest in the considered setting – distribution shifts without test time labels – is mainly motivated by the growing applications of CP in autonomous decision making. A very much incomplete list: see (Lekeufack et al., 2023) for a generic treatment; (Lindemann et al., 2023) for trajectory optimization in robotics; (Yang and Pavone, 2023; Gao et al., 2024) for 3D vision; (Kumar et al., 2023; Quach et al., 2024) for large language models (LLM); and (Ren et al., 2023) for LLM-powered robotics.
CP under distribution shifts
Considerable efforts have been devoted to developing CP methods robust to distribution shifts. We now discuss two possible directions and their respective limitations.
  • The first direction does not require test time labels, but the distribution shift is assumed to be simple in some sense. For example, Tibshirani et al. (2019) studied CP under covariate shifts, where the distribution of the label $Y$ conditioned on the covariate $X$ remains unchanged. Here, it suffices to use the classical likelihood ratio reweighting on the calibration dataset, but due to the difficulty of likelihood ratio estimation, one could also take a robust optimization perspective (Cauchois et al., 2024) and apply multicalibration on the possible test domain, as sketched in (Roth, 2022, Chapter 8). Barber et al. (2023) generalized the reweighting approach to handle mild general distribution shifts, but the choice of weights is generally unclear in practice.
  • The second direction is connecting CP to adversarial online learning. A line of works (Gibbs and Candes, 2021; Angelopoulos et al., 2023; Gibbs and Candès, 2023; Bhatnagar et al., 2023b; Zhang et al., 2024) applied regret minimization algorithms in OCO to select the score threshold in CP, and Bastani et al. (2022) achieved this task using multicalibration. By relaxing the CP objective from the coverage probability to the post-hoc coverage frequency, these methods can handle arbitrary continual distribution shifts. However, they require the true label to be provided after every prediction, which is a limiting requirement for many use cases. Our experiments show that it is possible to achieve comparable performance in these settings without this limitation, i.e., “label free”.
Test Time Adaptation (TTA)
@@ 2 Related works
Our EACP method is inspired by core ideas in TTA. The goal of TTA is to update a trained machine learning model at test time, using unlabeled data from shifted distributions. To achieve this, one could update the batch-norm statistics on the test data (Nado et al., 2020; Schneider et al., 2020; Khurana et al., 2021), or minimize the test-time prediction entropy (Wang et al., 2021; Zhang et al., 2022; Niu et al., 2022; Song et al., 2023; Press et al., 2024). Notably, these methods can be applied on any probabilistic and differentiable model (such as modern neural networks), which is naturally congruent with the strength of CP. However, to date this line of work has been disconnected from the conformal prediction literature, and to our knowledge there are no studies investigating the benefit of TTA to the distribution-shifted CP setting.
@@ 3 Preliminaries of CP
We begin by introducing the standard background of CP. For clarity, we assume i.i.d. data in our exposition, rather than the slightly weaker notion of exchangeability. Also see (Roth, 2022; Angelopoulos and Bates, 2023; Tibshirani, 2023).
@@
Let $\mathcal{D}$ be an unknown distribution on the space $\mathcal{X}\times\mathcal{Y}$ of covariate-label pairs, and let $\alpha\in(0,1)$ be the error rate we aim for. Given a dataset $D$ consisting of $n$ i.i.d. samples $(x_{i},y_{i})_{i\in[n]}\sim\mathcal{D}^{n}$, the goal of CP is to generate a set-valued function $\mathcal{C}_{D}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$, such that
$\mathbb{P}_{(x,y)\sim\mathcal{D},D\sim\mathcal{D}^{n}}\left[y\in\mathcal{C}_{D}(x)\right]\geq 1-\alpha.$ (1)
@@
That is, for a fresh sample $(x,y)\sim\mathcal{D}$, our prediction set $\mathcal{C}_{D}(x)$ covers the label $y$ with guaranteed high probability. [^3] In the literature this is called a marginal (w.r.t. the covariate $x$), in-expectation (w.r.t. the dataset $D$) guarantee – see (Roth, 2022, Chapter 7) for various ways to strengthen it. Notice that Eq.(1) alone is a trivial objective, since it suffices to predict the entire label space $\mathcal{C}_{D}(x)=\mathcal{Y}$ for all $x$. Therefore, CP is essentially a bi-objective problem: as long as Eq.(1) is satisfied, we want the prediction set $\mathcal{C}_{D}(x)$ to be small.
@@
The main difficulty of this set-valued prediction problem is that the range of output $2^{\mathcal{Y}}$ is too large. In this regard, the key idea of CP is reducing the problem to 1D prediction via a trained machine learning model, such as a neural network. Specifically, we assume access to a (positively oriented; i.e., larger is better) score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ given by the base model, such that for each covariate $x\in\mathcal{X}$ and candidate label $y^{\prime}\in\mathcal{Y}$, $s(x,y^{\prime})$ measures how likely the model believes that $y^{\prime}$ is the true label $y$. Then, all there is left for CP is to pick a threshold $\tau_{D}\in\mathbb{R}$ that depends on the dataset $D$, and predict the label set (if the score function is negatively oriented, then $\geq$ is replaced by $\leq$)
$\mathcal{C}_{D}\left(x;\tau_{D}\right):=\left\{y^{\prime}\in\mathcal{Y}:s\left(x,y^{\prime}\right)\geq\tau_{D}\right\}.$ (2)
@@
Under the i.i.d. assumption, the coverage objective Eq.(1) is satisfied by picking $\tau_{D}$ as the $\alpha(1+n^{-1})$-quantile of the empirical scores $\{s(x_{i},y_{i})\}_{i\in[n]}$. Since the training data of the base model is split from the calibration dataset used to determine $\tau_{D}$, this approach is commonly known as split conformal prediction, which we also refer to as SplitCP. Notably, $\tau_{D}$ is determined by the calibration dataset; once the latter is fixed, there is no need to access the ground truth labels at test time.
Example
In classification (which we focus on), the standard practice is to define the score function $s(x,y)=\pi_{\theta}(x)_{y}$ (Sadinle et al., 2019), where $\pi_{\theta}$ is a trained neural network parameterized by $\theta$, and $\pi_{\theta}(x)_{y}\in[0,1]$ is the softmax score corresponding to one of the $k$-classes $y\in[k]$. Such a score function is positively oriented (i.e., larger means more likely to be included in the prediction set).
@@
For the rest of this paper, we study a deviation of the above problem where the test sample $(x,y)$ is not drawn from $\mathcal{D}$, but rather some new unknown distribution $\widetilde{\mathcal{D}}_{\mathrm{test}}$. In general, it is impossible to guarantee Eq.(1) without restrictions on the shift, but we will show that with more help from the base model, the above CP procedure can be modified to work well in practice.

### v2

# Adapting Prediction Sets to Distribution Shifts Without Labels

@@ Abstract
Recently there has been a surge of interest to deploy confidence set predictions rather than point predictions in machine learning. Unfortunately, the effectiveness of such prediction sets is frequently impaired by distribution shifts in practice, and the challenge is often compounded by the lack of ground truth labels at test time. Focusing on a standard set-valued prediction framework called conformal prediction (CP), this paper studies how to improve its practical performance using only unlabeled data from the shifted test domain. This is achieved by two new methods called ECP and EACP, whose main idea is to adjust the score function in CP according to its base model’s own uncertainty evaluation. Through extensive experiments on a number of large-scale datasets and neural network architectures, we show that our methods provide consistent improvement over existing baselines and nearly match the performance of fully supervised methods. Our code can be found at: https://github.com/uoguelph-mlrg/EaCP.
footnotetext: Work done while visiting Harvard University.
@@ 1 Introduction
Advances in deep learning are fundamentally changing the autonomous decision making pipeline. While most works have focused on accurate point predictions, quantifying the uncertainty of the model is arguably as important. Taking autonomous driving for example: if a detection model predicts the existence of an obstacle, it would be reasonable to take different maneuvering strategies depending on the confidence of the prediction. But is that reliable? In a possible failure mode, the model could report $60\%$ (resp. $99\%$) confidence, but the probability of an obstacle actually showing up is $99\%$ (resp. $60\%$). Such discrepancy between the model’s own uncertainty evaluation and the ground truth probability (or post-hoc frequency) is commonly observed (Guo et al., 2017; Liang et al., 2023), and can compromise the safety in downstream decision making.
@@
Set-valued prediction provides an effective way to address this problem (Chzhen et al., 2021), with conformal prediction (CP; Vovk et al., 2005) being a well-known special case. Given a fixed black-box machine learning model (called the base model) and a covariate $x_{\mathrm{test}}$, the goal of CP is to generate a prediction set $\mathcal{C}_{\mathrm{test}}$ that contains (or, covers) the unknown ground truth label $y_{\mathrm{test}}$ with a pre-specified probability. Crucially, CP relies on the assumption that the distribution of the data stream is exchangeable (a weaker variant of i.i.d.), which allows the fairly straightforward inference of $y_{\mathrm{test}}$ from $x_{\mathrm{test}}$ and the base model’s performance on a pre-collected calibration dataset. Note that the ground truth label $y_{\mathrm{test}}$ does not need to be revealed after the set prediction is made: exchangeability together with a large enough labeled calibration dataset is sufficient to ensure the desirable coverage probability. This is particularly important for autonomous decision making, where real-time data annotation is expensive or even infeasible.
@@
However, real-world data streams are usually corrupted by all sorts of distribution shifts, violating the exchangeability assumption. Even when the data stream itself is exchangeable, we often want to continually update the base model rather than keeping it fixed, and this can be effectively understood as a distribution shift in the context of CP. In such cases, simply applying exchangeability-based methods could lead to highly inaccurate prediction sets (Tibshirani et al., 2019; Bhatnagar et al., 2023a; Kasa and Taylor, 2023). Therefore, making CP compatible with distribution shifts has become a focal point of recent works.
@@
A number of solutions have been proposed, but the key challenge still remains. For example, Gibbs and Candes (2021) formulated the connection between CP and Online Convex Optimization (OCO; Zinkevich, 2003), and the latter is able to handle arbitrarily distribution-shifted environments. The weakness is that ground truth labels are now required at test time (which we call full supervision), as opposed to the standard CP procedure. In the other direction, there are CP methods that combat distribution shifts without test time labels (Tibshirani et al., 2019; Barber et al., 2023; Cauchois et al., 2024), but they typically assume the distribution shifts are “easy”, such that even without labels, we can still rigorously infer the test distribution to a certain extent using the labeled calibration dataset. Overall, it appears that handling both difficulties – distribution shifts and the lack of test time labels – is a formidable but important remaining challenge.
@@
Focusing on classification, this paper develops practical unsupervised methods to improve the accuracy degradation of CP prediction sets under distribution shifts. The overarching idea is to exploit the uncertainty evaluation of the base model itself. Although such a quantity is not always calibrated in a strict sense, it has been consistently observed to strongly correlate with the magnitude of distribution shifts (Hendrycks and Gimpel, 2017; Wang et al., 2021; Kang et al., 2024), thus providing a valuable way to probe the test distribution without label access. Under this high level idea, we make the following contributions.
@@
  • First, we propose a new CP-inspired method named ECP (Entropy scaled Conformal Prediction). The key idea is to scale up the score function in standard CP by an “entropy quantile” of the base model, calculated on the unlabeled test dataset, which measures the base model’s own uncertainty on the test distribution.
@@
    More precisely, given each covariate $x_{\mathrm{test}}$ at test time, the score function in standard CP is determined by the fixed base model, and assigns each candidate label a “propensity score”. Then, the CP prediction set $\mathcal{C}_{\mathrm{test}}$ simply includes all the candidate labels whose score is above a certain threshold. [^1] The score functions are assumed to be positively oriented (Sadinle et al., 2019): labels with larger score are more likely to be included in the prediction set. By scaling up the score function while keeping the threshold fixed, ECP makes the prediction sets larger, which naturally corresponds to the intuition that the uncertainty of prediction should be inflated under distribution shifts. Moreover, the amount of such inflation is strongly correlated with the magnitude of the distribution shift, through the use of the entropy quantile.
@@
  • Second, we refine ECP using techniques from unsupervised Test Time Adaptation (TTA) (Niu et al., 2022), and the resulting method is named EACP (Entropy base-adapted Conformal Prediction). The key idea is that while ECP keeps the base model fixed at test time, we can concurrently update it using entropy minimization (Grandvalet and Bengio, 2004; Wang et al., 2021) – a widely adopted idea in unsupervised TTA, alongside the aforementioned entropy scaling. This “adaptively” reduces the scaling effect that ECP applies to the score function, thus shrinking the prediction sets of ECP smaller.
@@
  • Finally, we evaluate the proposed methods on a wide range of large-scale datasets under distribution shifts, as well as different neural network architectures. We find that exchangeability-based CP (with and without TTA on the base model) consistently leads to lower-than-specified coverage frequency. However, despite the absence of practical statistical guarantees in this setting, our methods can effectively mitigate this under-coverage issue while keeping the sizes of the prediction sets moderate. Furthermore, our methods also significantly improve the prediction sets generated by the base model itself (without CP). It shows that by bridging the CP procedure (which is statistically sound) and the base model’s own uncertainty evaluation (which is often informative), our methods enjoy the practical benefit from both worlds.
Related works
Considerable efforts have been devoted to developing CP methods robust to distribution shifts, which can be approximately categorized into two directions. The first direction does not require test time labels (Tibshirani et al., 2019; Cauchois et al., 2024), but the distribution shift is assumed to be simple in some sense. The second direction is connecting CP to adversarial online learning (Gibbs and Candes, 2021), but the true labels are required at test time. Due to space constraints, a thorough discussion is deferred to Appendix A, as well as a number of applications that motivate this work.
@@
Our techniques are inspired by core ideas in (unsupervised) TTA, whose goal is to update a trained machine learning model at test time, using unlabeled data from shifted distributions. To achieve this, one could update the batch-norm statistics on the test data (Nado et al., 2020; Schneider et al., 2020; Khurana et al., 2021), or minimize the test-time prediction entropy – a natural measure of the model’s uncertainty (Wang et al., 2021; Zhang et al., 2022; Niu et al., 2022; Song et al., 2023; Press et al., 2024). Notably, these methods can be applied to any probabilistic and differentiable model (such as modern neural networks), which is naturally congruent with the key strength of CP. However, to date this line of works has not been connected to the conformal prediction literature.
@@ 2 Preliminaries of CP
We begin by introducing the standard background of CP without distribution shifts. For clarity, we assume i.i.d. data in our exposition, rather than the slightly weaker notion of exchangeability. Also see (Roth, 2022; Angelopoulos and Bates, 2023; Tibshirani, 2023).
@@
Let $\mathcal{D}$ be an unknown distribution on the space $\mathcal{X}\times\mathcal{Y}$ of covariate-label pairs, and let $\alpha\in(0,1)$ be the error rate we aim for. Given a calibration dataset $D$ consisting of $n$ i.i.d. samples $\{x^{*}_{i},y^{*}_{i}\}_{i\in[n]}\sim\mathcal{D}^{n}$, the goal of CP is to generate a set-valued function $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$, such that
$\mathbb{P}_{(x_{\mathrm{test}},y_{\mathrm{test}})\sim\mathcal{D},D\sim\mathcal{D}^{n}}\left[y_{\mathrm{test}}\in\mathcal{C}(x_{\mathrm{test}})\right]\geq 1-\alpha.$ (1)
@@
That is, for a fresh test sample $(x_{\mathrm{test}},y_{\mathrm{test}})\sim\mathcal{D}$, our prediction set $\mathcal{C}(x_{\mathrm{test}})$ covers the ground truth label $y_{\mathrm{test}}$ with guaranteed high probability. Notice that Eq.(1) alone is a trivial objective, since it suffices to predict the entire label space $\mathcal{C}(x)=\mathcal{Y}$ for all $x$. Therefore, CP is essentially a bi-objective problem: as long as Eq.(1) is satisfied, we want the prediction set $\mathcal{C}(x)$ to be small.
@@
The main difficulty of this set-valued prediction problem is that the range of output $2^{\mathcal{Y}}$ is too large. In this regard, the key idea of CP is reducing the problem to 1D prediction via a trained machine learning model (called the base model), such as a neural network. Specifically, we assume access to a (positively oriented; i.e., larger is better) score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ given by the base model, such that for each test covariate $x_{\mathrm{test}}\in\mathcal{X}$ and candidate label $y\in\mathcal{Y}$, $s(x_{\mathrm{test}},y)$ measures how likely the model believes that $y$ is the true label $y_{\mathrm{test}}$. Then, all there is left for CP is to pick a threshold $\tau_{D}\in\mathbb{R}$ that depends on the dataset $D$, and predict the label set (if the score function is negatively oriented, then $\geq$ is replaced by $\leq$)
$\mathcal{C}(x_{\mathrm{test}}):=\left\{y\in\mathcal{Y}:s(x_{\mathrm{test}},y)\geq\tau_{D}\right\}.$ (2)
@@
Under the i.i.d. assumption, the coverage objective Eq.(1) is satisfied by picking $\tau_{D}$ as the $\alpha(1-n^{-1})$-quantile of the empirical scores $\{s(x^{*}_{i},y^{*}_{i})\}_{i\in[n]}$. Since the training data of the base model is split from the calibration dataset used to determine $\tau_{D}$, this approach is commonly known as split conformal prediction, which we refer to as SplitCP. Notably, $\tau_{D}$ is determined by the calibration dataset $D$; once the latter is fixed, there is no need to access the ground truth labels at test time.
Examples in classification
This paper focuses on classification. In this case, a simple and popular choice of the score function is $s(x,y)=\pi_{\theta}(x)_{y}$ (Sadinle et al., 2019), where $\pi_{\theta}$ is a trained neural network parameterized by $\theta$, and $\pi_{\theta}(x)_{y}\in[0,1]$ is the softmax score corresponding to one of the $k$-classes $y\in[k]$. Such a score function is positively oriented, which we adopt in this work. Another well-known choice due to Romano et al. (2020) is negatively oriented, and our methods can be applied there as well.
@@
For the rest of this paper, we study the following deviation of the above standard CP problem. At test time, instead of working with a single test sample $(x_{\mathrm{test}},y_{\mathrm{test}})$ drawn from $\mathcal{D}$, we consider a size-$N$ collection of samples [^2] The clearest notation is to index the test samples by $(x_{\mathrm{test},i},y_{\mathrm{test},i})$. Here we omit the subscript “$\mathrm{test}$” for conciseness. $\{x_{i},y_{i}\}_{i\in[N]}$ drawn from some new unknown distribution $\mathcal{D}_{\mathrm{test}}$. We only observe the covariates, defined as the test dataset $D_{\mathrm{test}}=\{x_{i}\}_{i\in[N]}$. Importantly, the ground truth labels on $D_{\mathrm{test}}$ are not revealed even after predictions are made. The goal, from a practical perspective, is to output a small prediction set $\mathcal{C}(x_{i})$ at each test covariate $x_{i}$, satisfying the specified empirical coverage rate,
$\frac{1}{N}\sum_{i=1}^{N}\bm{1}[y_{i}\in\mathcal{C}(x_{i})]\geq 1-\alpha.$
The function $\mathcal{C}$ can now depend on both the labeled calibration dataset $D$ and the unlabeled test dataset $D_{\mathrm{test}}$.
In general, it is impossible to prove meaningful bounds without assuming some form of similarity between $\mathcal{D}$ and $\mathcal{D}_{\mathrm{test}}$, but we will show that with help from the base model, the CP procedure can be modified to work well in practice.